OpenStudy (hafeda):

Suppose F(x)= k(2-x-x^3) has an inverse function and that f^-1 (3)= -2. Find k ?

OpenStudy (sunnnystrong):

\[f(x)=k(2-x-x^3)\] \[f^{-1}(3)=-2\] so... f(-2)=3 \[3=k(2-(-2)-(-2)^3)\] *Solve for k k=3/8

OpenStudy (hafeda):

Thanks a lot, i thought i had it wrong for sec. Thank u again